### v1

# Potential Field Based Deep Metric Learning

Abstract
Deep metric learning (DML) involves training a network to learn a semantically meaningful representation space. Many current approaches mine n-tuples of examples and model interactions within each tuplets. We present a novel, compositional DML model, inspired by electrostatic fields in physics that, instead of in tuples, represents the influence of each example (embedding) by a continuous potential field, and superposes the fields to obtain their combined global potential field. We use attractive/repulsive potential fields to represent interactions among embeddings from images of the same/different classes. Contrary to typical learning methods, where mutual influence of samples is proportional to their distance, we enforce reduction in such influence with distance, leading to a decaying field. We show that such decay helps improve performance on real world datasets with large intra-class variations and label noise. Like other proxy-based methods, we also use proxies to succinctly represent sub-populations of examples. We evaluate our method on three standard DML benchmarks- Cars-196, CUB-200-2011, and SOP datasets where it outperforms state-of-the-art baselines.
1 Introduction
The goal of visual metric learning is to learn a representation function where common distance metrics like the $\ell_{2}$ distance capture similarity between images. Semantic similarity is useful for applications such as image/video retrieval and search [3, 33, 21, 2], open set classification and segmentation [14, 6, 31], few-shot learning [39, 44], person re-identification[7, 53] and face verification [17, 30, 10]. Deep neural networks have helped make significant advances in visual metric learning [38], as they are trained to learn a non-linear mapping of images onto a semantic space.
Existing Methods: Popular loss functions to train such networks help force tuplets of examples/samples/embeddings of the same class to be closer than those in different classes. Examples include the triplet loss [51], the contrastive loss [8] and their extensions [40, 50].
The utility of supervision provided by samples in a tuplet depends on samples’ relative locations and labels [52]. Several strategies have been proposed to mine informative tuplets [38, 15], but they require computing the relative hardness of all tuples formed, which is computationally demanding (O($n^{3}$) for triplets among $n$ samples). Additionally, limiting interactions to tuplets incurs a loss of valuable supervision, greater susceptibility to noise [25], overfitting on the training set [26], and reduced learning of intra-class variations and generalizable features [36].
Proxy-based methods, first proposed in [33], use class representative ‘proxies’ as one or more hypothesized centers for each class. Computation of the loss is then in terms of sample-proxy distances instead of sample-sample distances. However, these methods learn less generalizable intra-class features due to the exclusive use of proxy-sample distances while ignoring sample-sample distances[37]. This effect is compounded when proxies, whose locations are themselves learned, end up far from sample embedding locations they are intended to represent.
Proposed Representation: To address these challenges, we propose Potential Field based Deep Metric learning (PFML), a metric learning framework that uses a continuous, potential field representation of data samples motivated by electrostatic fields used to model influence of charges in physics. Besides its analogical relevance, our motivation for using this representation comes from its previous, very successful uses, where it also helps overcome combinatorial complexity in modeling interactions. These applications include planning paths of a robot avoiding obstacles [19] and planning an efficient strategy for a robot manipulator to grasp an object [20].
In PFML, we view each data sample as a charge, creating a field. The fields due to the individual samples are added (superposed) to obtain the global potential field. This global field thus constitutes a compositional representation, capturing the combined influence at a given location (or sample) as a weighted addition of the influences of the individual samples. Each sample generates two types of fields: (1) an attraction field to drive similar examples closer to it, and (2) a repulsion field that drives dissimilar examples away from it, each weakening with distance.
We also make use of proxies like previous work, but only to augment the potential field generated by the current batch of samples. This is in contrast to past work that completely replaces sample-sample interactions with sample-proxy interactions. Preserving sample-sample relations through the potential field helps learn better intra-class features.
Advantages of Potential Field: (1) Using the potential field representation enables us to model interactions between all sample embeddings, as opposed to modeling those between small subsets (of sample or proxy points) as done in methods using e.g. point-tuplets based (e.g., contrastive) loss. (2) Addition of potentials due to all points also helps improve the quality of features learned while also (3) increasing robustness to noise since the effect of noise on interactions among a smaller number of samples will have a larger variance. (4) A major feature that differentiates our potential field is in the variation of strength of interaction between two points as the distance between them increases: instead of remaining constant or even becoming stronger, as is the case with most existing methods, e.g. [8, 21], in our model it becomes weaker with distance. [^1] It remains constant over a small radius and decays outside; however, for brevity, we will disregard this detail in the rest of this paper, and refer to the influence as decaying with distance, or as the decay property. The decay property is helpful in several ways: (4a) It ensures the intuitive expectation that two distant positive samples are too different to be considered as variants of each other, helping treat them as different varieties (e.g., associated with different proxies). (4b) The decay property also significantly improves PFML performance for the specific type of noise affecting labels, e.g., due to annotation errors common in real-world datasets (Sec. 4.3). (4c) As a result of the decay, the learned proxies remain closer to (at smaller Wasserstein distance $W_{2}$ from) the sample embeddings they represent than for (e.g., current proxy-based) methods where interactions strengthen with distance [21, 33, 35]; Sec 3.4), thereby enhancing their desired role.
Contributions: The main contributions of the proposed approach are:
  • We present PFML, a novel framework for metric learning with a compositional representation using a continuous, potential field to model all sample interactions instead of only interactions between n-tuplets. This enables better feature learning.
  • PFML reverses the almost universally used model of inter-sample interactions strengthening with distance; our decay property ensures that they weaken with distance. This (1) helps improve robustness in the presence of label noise, and (2) we prove that it helps proxies better align the distribution of proxies with the sample embeddings they represent.
  • We evaluate PFML on three zero-shot image retrieval benchmarks: Cars-196, CUB-200-2011, and the SOP dataset where it outperforms current, state-of-the-art techniques in both the standard no-label-noise scenario, and with even greater margins (>7% R@1) in the more realistic cases of label noise.
2 Related Work
Broadly, DML may be divided into two types, ranking/pair-based and proxy-based [21, 50].
Ranking/Pair based DML: These methods utilize losses that derive information from relations between inter-sample distances in pairs/ tuples of sample embeddings. Examples include the classical contrastive loss [8] or the triplet loss[51, 52, 38], which extends the concept to take into account relative distances in a triplet. [40] and [41] generalize these constraints to include multiple dissimilar samples. Mining informative pairs/tuples from a mini-batch plays an important role in the effectiveness of pair-based methods. Several strategies [38, 57, 15] have been proposed to pick samples based on their relative difficulty, with hard and semi-hard negative (picking pair of inter-class examples close together) and easy positive (same class examples close together) [54, 25] being popular. Our potential-field-based formulation naturally emphasizes such local (dis) similarities by virtue of its decaying formulation without the complexity of mining tuplets.
Proxy-based DML: ProxyNCA [33] and its variants [23, 46] attempt to minimize the distance of a sample from the proxy of its class while maximizing the distance from proxies of other classes. Proxy Anchor [21] builds on these by proposing a modified loss that takes into account sample hardness. [35, 10] introduced the use of multiple proxies per class to better model sub-clusters.
Potential Field: Potential fields have been used in a variety of applications to model interactions of objects. Of interest here is their use in robot path planning [19, 9], shape representation [18, 1] and object manipulation [20]. Here, they have been used to direct a robot towards its goal using an attraction field while avoiding obstacles using repulsion fields. Training supervision for DML shares a similar goal of moving sample embeddings towards other nearby samples of the same class, while avoiding dissimilar examples belonging to other classes.
3 Method
3.1 Problem Setup
The goal of deep metric learning is to define a semantic distance metric $d(\textbf{x}_{1},\textbf{x}_{2})$ for any pair of samples $\textbf{x}_{1},\textbf{x}_{2}\in\mathcal{D}$, where $\mathcal{D}$ =$=\{\textbf{x}_{i},y_{i}\},i\in\{1....\lvert\mathcal{D}\rvert\}$ and labels $y_{i}\in\{1,\ldots N\}$ represent N classes. The learned distance metric for any two points $d(\textbf{x}_{1},\textbf{x}_{2})$ should represent their semantic dissimilarity. Instead of directly learning the distance metric, a projector function $f_{\theta}$ is often represented by a neural network, parameterized by $\theta$ learned from data and distance $d(\textbf{x}_{1},\textbf{x}_{2})=\|f_{\theta}(\textbf{x}_{1}),f_{\theta}(\textbf{x}_{2})\|_{2}$ is given by the Euclidean distance between the projections/embeddings. For simplicity, we will refer to the embeddings $f_{\theta}(\textbf{x}_{i})$ belonging to a batch $\mathcal{B}$ sampled from $\mathcal{D}$ as $\textbf{z}_{i}$.
3.2 The Potential Field
With our potential field representation, embeddings need to be driven towards other nearby embeddings belonging to the same class, while also being driven away from embeddings of other classes. This is a mirror image of the behavior of an isolated system of electric charges, where dissimilar charges are drawn together while similar ones are repelled. We create attractive and repulsive electrostatic fields due to the given set of samples (charges) by summing up individual potentials exerted by each charge. The gradient of the net field on a sample yields the net electric force on the sample.
Figure 1: Overview of our Potential-field based DML pipeline. The process includes (1) Computing attraction and repulsion fields generated by each embedding and proxy, (2) Computing the class potential fields by superposition of individual fields (3) Evaluating total potential energy by summing up the potentials of embeddings and proxies under the class potential field and (4) Updating locations of sample embeddings (through network parameters) and proxies to minimize total potential energy through backprop.
Figure 2: An example of the class potential field $\Psi_{1}$ (Sec. 3.2.3), created by superposing the fields of individual embeddings belonging to Class 1. Arrows denote the gradient of the potential (force) which draws all class 1 embeddings together. Proxies at starred locations are drawn towards the nearest Class 1 embeddings which are potential minima.
Specifically, we define a potential field $\Psi_{j},$ for each class $\ j\in\{1,\ldots N\}$. The class potential field $\Psi_{j}$ is formed from a superposition of potentials belonging to sample embeddings from all classes. For an embedding from a given class, those other embeddings belonging to the same class have an attractive potential (to pull similar points closer), and those belonging to other classes have a repulsive potential (to push apart dissimilar points). We assume that each sample embedding $z_{i}$ has a unit charge, so the net force on it is the gradient of the potential field $\Psi_{y_{i}}$ of its class. We design the exact form of the potential functions by taking motivation from both, electrostatic fields and metric learning literature, which we discuss in detail in the following subsections. Figure 2 provides an overview of our complete pipeline of generation and use of the potential field for metric learning .
3.2.1 Field exerted by a single point
By using forces on the embeddings, we wish the training process to move the embeddings to a state of equilibrium at the end of training in which all embeddings from a class are nearby. To have a potential function exhibiting this behavior, we make use of the following two properties from electrostatics:
Property 1 - Equilibrium Property: Potential generated by an individual charge achieves an extremum at the embedding (minimum for attraction field, maximum for repulsion). This ensures that similar points are pulled together and dissimilar ones are pushed apart on minimizing potential.
Property 2 - Decay Property: The magnitude of the potential field due to a charge decays with distance. Such a decay ensures two desired characteristics. (1) An embedding is not pulled towards a distant sample embedding belonging to the same class, which is likely to be too different to be considered as a variant of it; instead, the decayed influence helps treat the distant sample as significantly different variety of the class, associated with other nearby embeddings of the class. (2) Like for a single charge in the Equilibrium property, a superposition of potentials of different embeddings still yields local minima at sample embeddings. As shown in Sec 3.4, this helps proxies come close to such minima and thus help better model the data distribution.
Each embedding $\textbf{z}_{i}$ generates an attraction potential $\psi_{att}$ and repulsion potential $\psi_{rep}$ whose strength at location r is given by:
$\psi_{att}(\textbf{r},\textbf{z}_{i}):$ =$=\begin{dcases}-\dfrac{1}{\delta^{\alpha}}&\textrm{if }\|\textbf{r}-\textbf{z}_{i}\|_{2}<\delta\\ -\dfrac{1}{\|\textbf{r}-\textbf{z}_{i}\|_{2}^{\alpha}}&\textrm{otherwise.}\end{dcases}$ (1)
$\psi_{rep}(\textbf{r},\textbf{z}_{i}):$ =$=\begin{dcases}\dfrac{1}{\|\textbf{r}-\textbf{z}_{i}\|_{2}^{\alpha}}&\|\textbf{r}-\textbf{z}_{i}\|_{2}<\delta\\ \dfrac{1}{\delta^{\alpha}}&\textrm{otherwise.}\end{dcases}$ (2)
Here, $\alpha$ is a hyperparameter that determines the rate of decay of the field (we evaluate it’s effect in Sec. 4.5 instead of fixing $\alpha=1$ as done in electrostatic fields). Note that the attraction potential is negative and increases with distance, implying that the force between the embedding and the point is attractive. The repulsion potential is positive and decreasing with distance, indicating repulsion.
A constant attraction potential inside the $\delta$ sphere (i.e., force = 0) ensures that sample embeddings within $\delta$ distance of one another do not continue to attract each other, thus preventing the network from embeddings from collapsing to a single point which harms generalizability of the features learned (analogous to the effect margins used in tuplet-based losses have).
Limiting the repulsion potential to a sphere of radius $\delta$ ensures that embeddings not belonging to the same class are driven to a distance $>\delta$. Since the aforementioned constant attraction potential ensures that $\delta$ is the maximum distance between samples $z_{i}$ from the same class, this leads to connected components of $\delta$ spheres formed by nearby embeddings of the same class. Extending repulsion to distances $>\delta$ does not provide any useful additional supervisory information while forcing further unmotivated inter-class separation; as we empirically show in Sec. 4.5, such excessive separation does indeed harm performance.
3.2.2 Proxies to Represent Sample Populations
To represent the image embeddings not sampled in the current batch $\mathcal{B}$, we propose to use a set of $M$ proxies per class for each of the N classes, $\textbf{p}_{j,k},k\in\{1,\ldots M\},j\in\{1,\ldots N\}$, and augment the potential field generated by the embeddings in the batch by those generated by the proxies. The proxies serve as stand-ins for the larger numbers of out-of-the-batch embeddings for calculating the attraction and repulsion fields, while also being affected by the field themselves like all embeddings. The proxies, being learnable parameters with high learning rates, tend to gravitate toward the nearest class modes of the potential field (and hence the class data distribution). The number of proxies M chosen should be large enough to capture the class-modes.
3.2.3 Class Potential Field
Given a batch $\mathcal{B}$ of samples and all proxies, we calculate the class potential field $\Psi_{j}(\textbf{r})$ (visualized in Fig 2) at a location r due to a class $j$ as a superposition of their attraction and repulsion potentials:
$\Psi_{j}(\textbf{r})$ $=\Psi_{j,att}(\textbf{r})+\Psi_{j,rep}(\textbf{r})$ (3)
where $\text{where }\Psi_{j,att}(\textbf{r})$ $=\sum_{i=1}^{\|\mathcal{B}\|,y_{i}=j}\psi_{att}(\textbf{r},\textbf{z}_{i})+\sum_{k=1}^{M}\psi_{att}(\textbf{r},\textbf{p}_{j,k})$ (4)
and $\text{and }\Psi_{j,rep}(\textbf{r})$ $=\sum_{i=1}^{\|\mathcal{B}\|,y_{i}\neq j}\psi_{rep}(\textbf{r},\textbf{z}_{i})+\sum_{\gamma=1,\gamma\neq j}^{N}\sum_{k=1}^{M}\psi_{rep}(\textbf{r},\textbf{p}_{\gamma,k})$ (5)
3.3 Training
Using gradient descent, the network is trained to minimize the total potential energy $\mathcal{U}$ of all proxies and the embeddings in the batch $\textbf{z}_{i}\in\mathcal{B}$ by applying a force on embeddings and proxies in the direction of $-\nabla\Psi_{y_{i}}(\textbf{z}_{i})$. $\mathcal{U}$ given by :
$\mathcal{U}=\sum_{i=1}^{\|\mathcal{B}\|}\Psi_{y_{i}}(\textbf{z}_{i})+\sum_{j=1}^{N}\sum_{k=1}^{M}\Psi_{j}(\textbf{p}_{j,k})$ (6)
3.4 Why use a decaying potential?
PFML uses a decaying potential, which yields two main benefits:
3.4.1 Better alignment of proxies with data distribution
The use of proxies in a framework where interactions decay with distance helps better align the distribution of proxies with the underlying sample data distribution (of the complete dataset) that they represent. We analyze this theoretically by constructing a potential field-based formulation where strength of interaction (defined in Eq. 1 & 2 for PFML) is directly proportional to the square of the distance from it ($\|\textbf{r}-\textbf{r}_{0}\|_{2}^{2}$) like it is in the original contrastive loss [8]. We keep all other details (use of $\delta$, proxies) as the same and call this the Contrastive Potential based Metric Learning (CPML), with its associated potential fields denoted by $\Psi_{att}^{*}\ ,\Psi_{rep}^{*}$ ( formally defined in Supplement Sec A). We analyze and compare CPML and PFML theoretically using the following proposition and its corollary:
Proposition 1:Let $Z=\{\textbf{z}_{1}\ldots\textbf{z}_{n}\}$ be a set of sample embeddings belonging to a class, then there exists a $0<\delta<\dfrac{\min_{i,j}\|z_{i}-z_{j}\|_{2}}{2\left(1+\frac{1}{n}\right)}$ and points $\textbf{z}_{min,i}$ within $\delta$ distance from each embedding $z_{i},\ i\in\{1\ldots n\}$ such that the attractive potential field $\Psi_{att}$ defined using $Z,\delta$ has a minimum at each $\textbf{z}_{min,i}$. The field $\Psi_{att}^{*}$ defined by them does not achieve a minimum at points within $\delta$ distance from all $z_{i}$.
Proposition 1 is a consequence of the potential field of an embedding dominating over other potential fields in its vicinity. Its proof relies on the Equilibrium and Decay properties, which are not satisfied by $\Psi^{*}$ and can be found in Supp. Sec A. A consequence of this is:
Corollary 1: Let $Z=\{\textbf{z}_{1}\ldots\textbf{z}_{n}\}$ be a set of sample embeddings from a class exerting an attraction field on a set of proxies $P=\{\textbf{ p}_{1}\ldots\textbf{p}_{m}\}$. Consider the equilibrium distribution $P_{eq}$ of proxies minimizing the potential energy. If the potential field is defined by $\Psi_{\text{att}}$, then the Wasserstein distance $W_{2}$ between $P_{eq}$ and the subset of data they represent is lower than when the potential field is defined by $\Psi_{\text{att}}^{*}$.
The proof for Corollary 1 relies on Proposition 1 and can be found in the Supplement (Sec. B).
Discussion: As proved in Corollary 1, CPML leads to proxies having a larger distance from the data distribution they represent (measured by $W_{2}$) than our proposed PFML, making the use of proxies in such a framework less effective. This leads to the proxies forming poorer stand-ins for the sample embeddings they represent, reducing the quality of features learned (verified in Sec. 4). This is caused by CPML not following the decay property (which is its only difference in design wrt PFML). Other proxy-based losses like the Proxy NCA [33], Soft Triplet [35], and Proxy Anchor [21] which perform similarly as CPML (see Tab. 1), also share with it the property of force increasing with distance (violating the decay property). Due to this, we expect them to similarly have a poorer alignment of proxies with underlying data distribution as compared to PFML (Applying an analysis similar to the one in Proposition 1 to them is considerably more challenging due to the absence of a potential field structure.). We verify this empirically in Sec. 4.4 by measuring the average distance between proxies and the closest data points ($W_{2}$ distance) for these methods.
3.4.2 Robustness when learning with noisy labels
The decay property helps PFML become significantly more robust to the effect of incorrectly labeled examples as compared to existing methods. Recent work by [26] shows that such examples are common in metric learning benchmarks. The decay de-emphasizes the contribution of such mislabeled examples (typically located away from correctly labeled samples). We empirically confirm this in Sec. 4.3, demonstrating that methods not adhering to the decay property (i.e., all previous tuplet-based and proxy-based methods) experience significantly greater performance degradation due to label noise compared to PFML.
3.5 Advantages in Learning Fine-grained Features
In contrast with all previous proxy-based methods, PFML (1) Models sample-sample interactions directly using the potential field, gaining fine-grained supervision, and (2) Makes use of an explicit margin-like $\delta$ for proxies, which has been shown to help zero-shot generalization and fine-grained feature learning in tuplet-based methods[52]. Turning off the interaction within the $\delta$ radius also prevents a bias towards learning isotropic distributions as is present in other proxy-based methods [37], helping us obtain better intra-class resolution.
4 Experiments and Results
Benchmarks $\rightarrow$ CUB-200-2011 Cars-196 SOP Methods $\downarrow$ (Chronological) R@1 R@2 R@4 R@1 R@2 R@4 R@1 R@10 R@100 ResNet50 (512 dim) ESPHN [54] 64.9 75.3 83.5 82.7 89.3 93.0 78.3 90.7 96.3 N.Softmax [58] 61.3 73.9 83.5 84.2 90.4 94.4 78.2 90.6 96.2 DiVA [32] 69.2 79.3 - 87.6 92.9 - 79.6 91.2 - Proxy NCA++ [46] 64.7 - - - 85.1 - - 79.6 - - Proxy Anchor [21] 69.7 80.0 87.0 87.7 92.9 95.8 - - - DCML-MDW [59] 68.4 77.9 86.1 85.2 91.8 96.0 79.8 90.8 95.8 MS+DAS [29] 69.2 79.2 87.1 87.8 93.1 96.0 80.6 91.8 96.7 HIST [28] 71.4 81.1 88.1 89.6 93.9 96.4 81.4 92.0 96.7 HIER[22] 70.1 79.4 86.9 88.2 93.0 95.6 80.2 91.5 96.6 HSE-PA [55] 70.6 80.1 87.1 89.6 93.8 96.0 80.0 91.4 96.3 CPML (Sec. 3.4) 68.3 78.7 86.2 85.2 91.5 95.2 79.4 90.7 96.1 Potential Field (Ours) 73.4 $\pm$ 0.3 82.4$\pm$ 0.1 88.8$\pm$ 0.1 92.7$\pm$ 0.3 95.5 $\pm$ 0.1 97.6$\pm$ 0.1 82.9 $\pm$ 0.2 92.5 $\pm$ 0.2 96.8 $\pm$ 0.1 BN Inception (512 dim) HTL [13] 57.1 68.8 78.7 81.4 88.0 92.7 74.8 88.3 94.8 MultiSimilarity [50] 65.7 77.0 86.3 84.1 90.4 94.0 78.2 90.5 96.0 SoftTriple [35] 65.4 76.4 84.5 84.5 90.7 94.5 78.3 90.3 95.9 CircleLoss [43] 66.7 77.4 86.2 83.4 89.8 94.1 78.3 90.5 96.1 DiVA [32] 66.8 77.7 - 84.1 90.7 - 78.1 90.6 - ProxyGML [61] 66.6 77.6 86.4 85.5 91.8 95.3 78.0 90.6 96.2 Proxy Anchor [21] 68.4 79.2 86.8 86.1 91.7 95.0 79.1 90.8 96.2 DRML-PA [60] 68.7 78.6 86.3 86.9 92.1 95.2 71.5 85.2 93.0 MS+DAS [29] 67.1 78.11 86.4 85.7 91.6 95.3 78.2 90.3 96.0 HIST [28] 69.7 80.0 87.3 87.4 92.5 95.4 79.6 91.0 96.2 DFML-PA [48] 69.3 - - 88.4 - - - - - HSE-M [55] 67.6 78.0 85.8 82.0 88.9 93.3 - - - PA+niV [23] 69.5 80.0 - 86.4 92.0 - 79.2 90.4 - Potential Field (Ours) 71.5$\pm$ 0.3 81.2$\pm$0.2 88.3$\pm$0.2 90.1$\pm$0.2 93.9$\pm$0.1 96.3$\pm$0.1 80.6 $\pm$ 0.3 91.8 $\pm$ 0.1 96.4 $\pm$ 0.1 DINO (384 dim) DINO [5] 70.8 81.1 88.8 42.9 53.9 64.2 63.4 78.1 88.3 Hyp [12] 80.9 87.6 92.4 89.2 94.1 96.7 85.1 94.4 97.8 HIER [22] 81.1 88.2 93.3 91.3 95.2 97.1 85.7 94.6 97.8 Potential Field (Ours) 83.1 $\pm$ 0.3 89.3 $\pm$ 0.2 94.2 $\pm$ 0.1 94.7 $\pm$ 0.1 96.5 $\pm$ 0.1 97.8 $\pm$ 0.1 86.5 $\pm$ 0.3 95.1 $\pm$ 0.3 98.0 $\pm$ 0.2 ViT (384 dim) ViT-S [11] 83.1 90.4 94.4 47.8 60.2 72.2 62.1 77.7 89.0 Hyp [12] 85.6 91.4 94.8 86.5 92.1 95.3 85.9 94.9 98.1 HIER [22] 85.7 91.3 94.4 88.3 93.2 96.1 86.1 95.0 98.0 Potential Field (Ours) 87.8 $\pm$ 0.2 92.6 $\pm$ 0.2 95.7 $\pm$ 0.1 91.5 $\pm$ 0.3 95.2 $\pm$ 0.2 97.4 $\pm$ 0.1 88.2 $\pm$ 0.1 95.7 $\pm$ 0.1 98.6 $\pm$ 0.1
[…]
4.1 Setup
Datasets: We empirically compare PFML with current state-of-the-art baselines on three public benchmark datasets (1) The Cars-196 dataset [24] which includes 16,185 images from 196 car model categories, (2) CUB-200-2011 dataset [47] consisting of 11,788 images from 200 bird species, and (3) Stanford Online Products (SOP) dataset [41] containing 120,053 images of 22,634 different products sold online. The training and testing splits for the three datasets follow standard zero-shot retrieval-based evaluation protocol [50, 21, 61] of using the first half of the classes for training and the second half for testing.
Backbone: To enable standardized comparison with a wide variety of methods, we evaluate our method using both the Inception with batch normalization (IBN) [45], ResNet-50 [16], ViT [11] and DINO [5] backbones commonly used by DML methods. ImageNet pre-training is used for initialization, and only the last fully connected layer is changed to fit the dimension of the embedding needed. We follow past work [21, 46, 35] and $\ell_{2}$ normalize the final output. An input image size of $224\times 224$ is used for all experiments and comparisons.
Training parameters: We train the backbone for 200 epochs on each dataset. We use the Adam optimizer with the learning rate chosen as $5e^{-4}$. We scale up the learning rate of our proxies by 100 times following [21, 33, 35] for better convergence. We choose values of $\delta$ between $[0.1,0.3]$, $\alpha$ between $\{0,6\}$ via cross-validation. We use $M=15$ for the CUB-200 and Cars-196 datasets, while $M=2$ is used for the SOP dataset. All experiments are performed on a single NVIDIA A4000 GPU.
Evaluation Settings: We follow [21, 29, 50, 22] in using $224\times 224$ sized center crops from $256\times 256$ images for evaluation. We measure image retrieval performance using Recall@K, which computes the percentage of samples which have a valid similar neighbor (belonging to same class) among its K nearest neighbors.
4.2 Image Retrieval Performance
As seen in Table 1, our method outperforms all other methods on Image Retrieval on the CUB-200-2011, Cars-196 and SOP datasets as measured by the Recall@K metric. It consistently outperforms other methods irrespective of the backbone used, significantly outperforming the best-performing proxy-based methods (which similar to us, only change the loss function) like ProxyAnchor[21], ProxyGML[61] and SoftTriplet [35] in terms of Recall@1 (R@1), by more than $5\%$ on Cars-196, $3.7\%$ on CUB-200-2011 and $1.5\%$ on the SOP dataset. It also outperforms the significantly more complex, current state-of-the-art HIST[28] by $3.1\%$, $2\%$ and $1.5\%$ in terms of R@1 when using the Resnet-50 backbone on the Cars-196, CUB-200-2011 and SOP datasets respectively. It also significantly outperforms the best tuplet-based loss: the Multi-Similarity loss by $6\%$, $5.8\%$ and $2.4\%$ in terms of R@1 when using the BN Inception backbone on the Cars-196, CUB-200 and SOP datasets respectively. Note that our performance gains are larger than current SOTA compared to improvements reported by previous efforts on the same benchmarks, despite our intuitive and simpler method.
We also evaluate our method 1) Using the protocol and metrics proposed in [34] and (2) When used to learn smaller 128-dim embeddings for more efficient retrieval. Our method still outperforms all state-of-the-art methods in both cases. Details of these experiments are in Supp. Sec. C.
4.3 Image Retrieval in Presence of Label Noise
[…]
Figure 3: Variation in Recall@1 with $M$, $\delta$, $\alpha$ and $\delta_{rep}-\delta_{att}$ on the Cars-196 and CUB-200-2011 datasets. Error bars represent std. deviations over 5 runs.
We evaluate the robustness of our method vs. state-of-the-art baselines in the realistic scenario of label noise (e.g., due to mislabeling, which is very common, particularly in fine-grained datasets such as those used for DML). Specifically, we randomly corrupt labels of 20% of samples in the training set. As seen in the results in Table 3, this leads to a significant decline in the performance of all methods. But our method is the least affected by the noise. Other methods like the triplet [15] or ProxyAnchor[21] suffer much greater declines because they are affected more by distant points represented by the mislabeled examples. Our method significantly outperforms Proxy Anchor [21] (the second-best method) by $6\%$ and $7.6\%$ in terms of R@1 on the CUB-200-2011 and Cars-196 datasets.
4.4 Distance Between Proxies and Sample Embeddings
We measure the average (over a complete epoch) Wasserstein distance $W_{2}$ between proxies and the subset of training data represented by/closest to them  [^2] For methods using a single proxy per class, this equals the distance between a proxy and the closest embedding belonging to its class for different proxy-based methods. We train networks on the CUB-200-2011 using the settings in Section 4.1 and use the same number of proxies for all methods utilizing multiple proxies. As seen in Table 3, our method shows 33.3% lower $W_{2}$ distance between the proxies and the subset of data than the second best SoftTriplet [35] loss, empirically validating our discussion in Sec. 3.4.1.
4.5 Ablation studies
In all experiments, we use the parameters mentioned in Section 4.1 with a ResNet-50 backbone on both the Cars-196 and CUB-200-2011 datasets.
4.5.1 Effect of proxies and their number
Proxies augment the potential field, and the number of proxies per class $M$ decides the granularity at which the proxies model the distribution of sample embeddings. We study their effect on performance by varying $M$ between $[0,30]$. From the results plotted in Figure 3(a) we observe that (1) Using proxies to augment the field boosts performance (vs using no-proxies) and (2) Network performance remains stable for a wide range of M, decreasing slightly for small values of M (still outperforming most other proxy-based methods). This behavior is expected as using potential field created using a very small M would not be able to effectively model underlying sample distribution, leading to worse supervision. Additionally, we observe that using a larger number of proxies per class does not significantly affect the training times, as bulk of computation is dominated by the network.
4.5.2 Effect of varying $\delta$
We conduct an empirical study to determine the effect of varying the radius parameter $\delta$ on performance by varying it over $[0.1,0.35]$. As seen in Figure 3(b), we observe that performance remains relatively stable for a wide range of $\delta$ values, demonstrating our method’s robustness to it.
4.5.3 Effect of potential decay $\alpha$.
The decay parameter $\alpha$ determines the rate at which the potential field of an embedding decays with distance. We study its effect by varying it over $\left[0,6\right]$. As seen in Figure 3(c), performance is worse when the potential field does not decay ($\alpha=0$). A field that decays strongly $\alpha\in[3,6]$ seems to perform better than one with a much milder decay ($\alpha\in\left[1,2\right]$). The likely reason for this trend is that when the potential field has no decay /decays very mildly, the force on an embedding would still be dominated by a large number of distant sample embeddings rather than those in its vicinity.
4.5.4 Restricting repulsion to a distance of $\delta$
In contrast with most proxy-based methods, our design of potential field $\Psi_{rep}$ limits the range of repulsion to within $\delta$ distance from embeddings in proxies. We study the effect of extending the range of repulsion beyond the $\delta$ radius by having $\delta_{rep}>\delta_{att}$ where $\delta_{att}$ and $\delta_{rep}$ represent the values of $\delta$ used for attraction (Eq.1) and repulsion (Eq.2) fields, respectively. We fix $\delta_{att}=0.2$ for this experiment, increasing $\delta_{rep}$ to increase their difference. As observed in Figure 3(d), we see that this leads to a decline in performance, validating our design. This is due to individual potentials repelling dissimilar points to a distance of at least $\delta_{rep}$, greater than the distance $\delta_{att}$ within which all points of the same class are attracted to. Using $\delta_{rep}>\delta_{att}$ enforces excessive separation between dissimilar examples, compressing intra-class features more than needed.
5 Limitations
PFML relies on the use of proxies to augment potential field of samples within a batch. Hence, it shares the main limitation of proxy-based methods, i.e., when the number of classes is large, proxies need a correspondingly large amount of compute during training because the number of proxies scales linearly with the number of classes. Despite this limitation, proxy-based methods have been widely used in applications with a large number of classes like face recognition (e.g. [10] on 500,000 classes), likely due to (1) Use of proxies only affects training, not inference complexity and (2) The parameters/compute complexity added by the proxies is relatively low for the performance gains they offer in several applications, especially when such compute complexity is compared to the forward and backward passes of the neural network that consume bulk of the compute in DML.
6 Conclusion
In conclusion, PFML introduces a novel paradigm for capturing the interactions between all samples (as opposed to tuples) using a continuous potential field. In our potential field, the strength of sample influence decreases with distance, reversing the common model of increasing influence with distance. This helps significantly improve robustness under real-world like label noise, while also helping better align the distribution of proxies with the distribution of the sample embeddings they represent. The overall design of our method is validated by its performance on standard image retrieval benchmarks, where it outperforms state-of-the-art methods both in the standard noise-free setting and in the presence of label noise unavoidable in real-world datasets.
Appendix A Proof of Proposition 1
Proposition 1:Let $Z=\{\textbf{z}_{1}\ldots\textbf{z}_{n}\}$ be a set of sample embeddings belonging to a class, then there exists a $0<\delta<\dfrac{\min_{i,j}\|z_{i}-z_{j}\|_{2}}{2\left(1+\frac{1}{n}\right)}$ and points $\textbf{z}_{min,i}$ within $\delta$ distance from each embedding $z_{i},\ i\in\{1\ldots n\}$ such that the attractive potential field $\Psi_{att}$ (Eq. 4) defined using $(Z,\ \delta)$ has a minimum at each $\textbf{z}_{min,i}$. The field $\Psi_{att}^{*}$ defined by them does not achieve a minimum at points within $\delta$ distance from all $z_{i}$.
Proof: We first define the potential fields of an individual embedding $\textbf{z}_{i}$, $\psi_{att}(\textbf{r},\textbf{z}_{i})$ and $\psi_{att}^{*}(\textbf{r},\textbf{z}_{i})$ are defined as:
$\psi_{att}(\textbf{r},\textbf{r}_{0}):=\begin{dcases}-\dfrac{1}{\delta^{\alpha}}&\textrm{if }\|\textbf{r}-\textbf{r}_{0}\|_{2}<\delta\\ -\dfrac{1}{\|\textbf{r}-\textbf{r}_{0}\|_{2}^{\alpha}}&\textrm{otherwise.}\end{dcases}$
$\psi_{att}^{*}(\textbf{r},\textbf{r}_{0}):=\begin{dcases}\delta^{2}&\textrm{if }\|\textbf{r}-\textbf{r}_{0}\|_{2}<\delta\\ \|\textbf{r}-\textbf{r}_{0}\|_{2}^{2}&\textrm{otherwise.}\end{dcases}$
The potential fields created by all data points are:
$\Psi_{att}(\textbf{r})=\sum_{i=1}^{n}\psi_{att}(\textbf{r},\textbf{z}_{i})$ (7)
$\Psi_{att}^{*}(\textbf{r})=\sum_{i=1}^{n}\psi_{att}^{*}(\textbf{r},\textbf{z}_{i})$ (8)
We prove the proposition in two parts, first proving the assertion for $\Psi_{att}(\textbf{r})$ in part 1, and then moving on to proving the assertion for $\Psi_{att}^{*}(\textbf{r})$ in part 2.
A.1 Part 1: Proof for $\Psi_{att}(\textbf{r})$
$\Psi_{att}$ is continuous and bounded within the hyperspheres $S_{i}=\|\textbf{r}-\textbf{z}_{i}\|\leq\delta$. The hyperspheres form a closed compact set. By Bolzano Weierstrass theorem, $\Psi_{att}$ achieves a minimum $\Psi_{att}(\textbf{z}_{min,i})$ on the set $S_{i}$ at $\textbf{z}_{min,i}$. $\textbf{z}_{min,i}$ may lie either inside the sphere or on the boundary. We analyze both cases separately separately:
Case 1: If $\textbf{z}_{min,i}$ lies inside the sphere $S_{i}$, then $\textbf{z}_{min,i}$ is a minimum for $\Psi_{att}$ on $R^{D}$ too because $S_{i}\subset R^{D}$ (D = embedding dimension).
Case 2: If $\textbf{z}_{min,i}$ lies on the border of sphere $S_{i}$, then by the definition of a local minima
$\textbf{z}_{min,i}$ is a minimum of $\Psi_{att}$ on $R^{D}$ iff there exists an $\epsilon>0$ such that $\Psi_{att}(\textbf{z}_{min,i})\leq\Psi_{att}(\textbf{z}^{{}^{\prime}})$ for all $\|\textbf{z}^{{}^{\prime}}-\textbf{z}_{min,i}\|<\epsilon$.
For all $\textbf{z}^{{}^{\prime}}\in S_{i}$, we know that $\Psi_{att}(\textbf{z}_{min,i})\leq\Psi_{att}(\textbf{z}^{{}^{\prime}})$ by definition of $\textbf{z}_{min,i}$. If $\textbf{z}^{{}^{\prime}}\notin S_{i}$, then using Taylor expansion at $\textbf{z}_{min,i}$,
$\Psi_{att}(\textbf{z}^{{}^{\prime}})=\Psi_{att}(\textbf{z}_{min,i})+(\textbf{z}^{{}^{\prime}}-\textbf{z}_{min,i}).\nabla\Psi_{att}(\textbf{z}_{min,i})+\text{err}(\textbf{z}^{{}^{\prime}}-\textbf{z}_{min,i})$ (9)
Here, $\text{err}(\textbf{z}^{{}^{\prime}}-\textbf{z}_{min,i})$ consists of all higher degree terms of the Taylor expansion.
Now, the term $\nabla\Psi_{att}(\textbf{z}_{min,i})$ in the taylor expansion is given by:
$\nabla\Psi_{att}(\textbf{z}_{min,i})$ $=\dfrac{\alpha}{\|\textbf{z}_{min,i}-\textbf{z}_{i}\|_{2}^{\alpha+1}}\hat{\mathbf{r}}+\gamma_{i}$
$\nabla\Psi_{att}(\textbf{z}_{min,i})$ $=\dfrac{\alpha}{\delta^{\alpha+1}}\hat{\mathbf{r}}+\gamma_{i}\hat{\mathbf{r}}$
where $\text{where }\gamma_{i}$ $=\sum_{j=1,\neq i}^{n}\nabla\psi_{att}(\textbf{z}_{min,i}-\textbf{z}_{j})$
Now $\dfrac{\alpha}{\delta^{\alpha+1}}>\gamma_{i}$ for all $\delta<\left(\dfrac{\alpha}{\gamma_{i}}\right)^{\frac{1}{\alpha+1}}$. So,
$(\textbf{z}^{{}^{\prime}}-\textbf{z}_{min,i}).\nabla\Psi_{att}(\textbf{z}_{min,i})=\|\textbf{z}^{{}^{\prime}}-\textbf{z}_{min,i}\|\left(\dfrac{\alpha}{\delta^{\alpha+1}}+\gamma_{i}\right)>0\ \ \ \forall\delta<\left(\dfrac{\alpha}{\gamma_{i}}\right)^{\frac{1}{\alpha+1}}$ (10)
Also, as $\epsilon\to 0$, $\dfrac{\text{err}(\textbf{z}^{{}^{\prime}}-\textbf{z}_{min,i})}{(\textbf{z}^{{}^{\prime}}-\textbf{z}_{min,i}).\nabla\Psi_{att}(\textbf{z}_{min,i})}\to 0$ as those are higher order terms.
Simplifying the Taylor expansion (Eq. 9)
$\Psi_{att}(\textbf{z}^{{}^{\prime}})-\Psi_{att}(\textbf{z}_{min,i})$ $=(\textbf{z}^{{}^{\prime}}-\textbf{z}_{min,i}).\nabla\Psi_{att}(\textbf{z}_{min,i})+\text{err}(\textbf{z}^{{}^{\prime}}-\textbf{z}_{min,i})=(\textbf{z}^{{}^{\prime}}-\textbf{z}_{min,i}).\nabla\Psi_{att}(\textbf{z}_{min,i})\text{ as }\epsilon\to 0$
[…]
So for $\textbf{z}^{{}^{\prime}}\notin S_{i}$, $\Psi_{att}(\textbf{z}_{min,i})\leq\Psi_{att}(\textbf{z}^{{}^{\prime}})$ for all $\|\textbf{z}^{{}^{\prime}}-\textbf{z}_{min,i}\|<\epsilon$.
Therefore, by definition $\textbf{z}_{min,i}$ is a minimum of $\Psi_{att}$ on $R^{D}$ in Case 2 too.
Hence, $\Psi_{att}$ achieves a minimum at $z_{min,i}\forall i\in\{1\ldots n\}$ for $\delta<\operatorname*{arg\,min}_{i}\left(\dfrac{\alpha}{\gamma_{i}}\right)^{\frac{1}{\alpha+1}}$. $z_{min,i}$ lies within $\delta$ distance of the embeddings $z_{i}$ by definition of the hyperspheres $S_{i}$.
A.2 Part 2: Proof for $\Psi_{att}^{*}(\textbf{r})$
Outside the hyperspheres $S_{i}=\|\textbf{r}-\textbf{z}_{i}\|\leq\delta$, $\nabla\Psi_{att}^{*}(\textbf{r})$ ($\textbf{r}\notin S_{i}$)is given by:
$\nabla\Psi_{att}^{*}(\textbf{r})$ $=\sum_{i=1}^{n}\nabla\psi_{att}^{*}(\textbf{r}-\textbf{z}_{i})$
$\nabla\Psi_{att}^{*}(\textbf{r})$ $=\sum_{i=1}^{n}2(\textbf{r}-\textbf{z}_{i})=2n\textbf{r}-\sum_{i=1}^{n}2\textbf{z}_{i}=2n\left(\textbf{r}-\sum_{i=1}^{n}\dfrac{\textbf{z}_{i}}{n}\right)$ (12)
It achieves a single minimum at $\textbf{z}_{globalmin}=\dfrac{\sum_{i=1}^{n}\textbf{z}_{i}}{n}$.
Now define differentiable extensions of the potentials-
$\psi_{att}^{**}(\textbf{r}-\textbf{r}_{0})=\|\textbf{r}-\textbf{r}_{0}\|_{2}^{2}$ (13)
To calculate $\nabla\Psi_{att}^{*}(\textbf{r})$ inside the hyperspheres ($\textbf{r}\in S_{i}$), we note that no two hyper spheres $S_{i}$ intersect with each other as $\delta<0.5\min_{i,j}\|z_{i}-z_{j}\|_{2}$. Observing that for $\textbf{r}\in S_{i}$
$\nabla\Psi_{att}^{*}(\textbf{r})=\sum_{j=1,j\neq i}^{n}\nabla\psi_{att}^{**}(\textbf{r}-\textbf{z}_{j})$
Using triangle inequality for $\textbf{r}\in S_{i}$:
$\|\sum_{j=1}^{n}\nabla\psi_{att}^{**}$ $(\textbf{r}-\textbf{z}_{j})\|-\|\nabla\psi_{att}^{**}(\textbf{r}-\textbf{z}_{i})\|\leq\|\sum_{j=1,j\neq i}^{n}\nabla\psi_{att}^{*}(\textbf{r}-\textbf{z}_{j})\|=\|\nabla\Psi_{att}^{*}(\textbf{r})\|$
[…]
$\|\nabla\Psi_{att}^{*}(\textbf{r})\|$ $\geq\|\sum_{j=1}^{n}\nabla\psi_{att}^{**}(\textbf{r}-\textbf{z}_{j})\|-2\delta$
We know that the RHS term $>0$ because $\|\sum_{j=1}^{n}\nabla\psi_{att}^{**}(\textbf{r}-\textbf{z}_{j})\|>2\delta$ for
$\|\textbf{r}-\sum_{i=1}^{n}\dfrac{\textbf{z}_{i}}{n}\|>\dfrac{\delta}{n}$ (14)
using Equation 12. For all such r:
$\|\nabla\Psi_{att}^{*}(\textbf{r})\|$ $>0$ (15)
At most one sphere $S_{i}$ has an $\textbf{r}_{i}\in S_{i}$ not satisfying Equation 14. We prove this by contradiction. Assume that another sphere $S_{j}$ has $\textbf{r}_{j}\in S_{j}$ not satisfying Equation 14. Now the distance between $\textbf{r}_{i}$ and $\textbf{r}_{j}$ satisfies:
$\|\textbf{r}_{i}-\textbf{r}_{j}\|\geq d_{min}-2\delta$ (16)
here $d_{min}=\min_{i,j}\|z_{i}-z_{j}\|_{2}$ Using the fact that $\delta<\dfrac{d_{min}}{2\left(1+\frac{1}{n}\right)}$ by definition and substituing for $d_{min}$ in equation 16 we get:
$\|\textbf{r}_{i}-\textbf{r}_{j}\|>2\delta\left(1+\dfrac{1}{n}\right)-2\delta$
$\|\textbf{r}_{i}-\textbf{r}_{j}\|>\dfrac{2\delta}{n}$
[…]
Hence for any other hypersphere $S_{j},i\neq j$, for all r we have $\|\nabla\Psi_{att}^{*}(\textbf{r})\|>0$, and hence no minimum exists within them. Hence, proved.
Appendix B Proof of Corollary 1
Corollary 1: Let $Z=\{\textbf{z}_{1}\ldots\textbf{z}_{n}\}$ be a set of sample embeddings belonging to a class exerting an attraction field on a set of proxies $P=\{\textbf{ p}_{1}\ldots\textbf{p}_{m}\}$. Consider the equilibrium distribution $P_{eq}$ of proxies minimizing the potential energy. If the potential field is defined by $\Psi_{\text{att}}$, then the Wasserstein distance $W_{2}$ between $P_{eq}$ and the subset of data they represent is lower than when the potential field is defined by $\Psi_{\text{att}}^{*}$.
Proof: Let the potential fields $\Psi_{\text{att}}$ and $\Psi_{\text{att}}^{*}$ be given by Equations 7 and 8 respectively.
$\psi_{total}(\textbf{r})=\sum_{i=1}^{n}\psi(\textbf{r}-\textbf{z}_{i})$ (17)
where $\psi$ is defined using 4 (att and class subscript j omitted for clarity). The potential energies of the proxy distribution $P_{eq}$ in these potential fields, $\mathcal{U}_{proxy}$ and $\mathcal{U}_{proxy}^{*}$ respectively are given by :
$\mathcal{U}_{proxy}=\sum_{p_{i}\in P_{eq}}\Psi_{\text{att}}(\textbf{p}_{i})$
$\mathcal{U}_{proxy}^{*}=\sum_{p_{i}\in P_{eq}}\Psi_{\text{att}}^{*}(\textbf{p}_{i})$
At equilibrium, each proxy migrates to the nearest minimum in the field. The subset of data the proxies represent are given by the subset of m data points $Z_{subset}=z_{f(k)},\ k\in\{1\ldots m\};Z_{subset}\subset Z$ which is the closest in Wasserstein distance $W_{2}$ from the proxies.
Case 1: Let the field be defined by $\Psi_{\text{att}}$. The proxy $\textbf{p}_{k},k\in{1\ldots m}$ migrates to the nearest stable minimum in the potential field denoted by $z_{min,f(k)}$ which is located within $\delta$ distance of data point $z_{f(k)}$ ( using proposition 1 ). Assuming that the proxies are initialized using a normal distribution (commonly used) and $m<<n$, which is typically true, we ignore the probability of more than one proxy going to the same minimum $z_{min,f(k)}$ (so $f(k)$ is one-one). Therefore:
$W_{2}(P_{eq},Z_{subset})$ $=\inf_{\pi}\left(\dfrac{1}{m}\sum_{k=1}^{m}\|p_{k}-z_{\pi(k)}\|_{2}\right)$
here the infimum is over all permutations $\pi$ of k elements
$W_{2}(P_{eq},Z_{subset})$ $=\left(\dfrac{1}{m}\sum_{k=1}^{m}\|z_{min,f(k)}-z_{f(k)}\|_{2}\right)$
Using proposition 1
$W_{2}(P_{eq},Z_{subset})$ $\leq\left(\dfrac{1}{m}\times(m\delta)\right)$
$W_{2}(P_{eq},Z_{subset})$ $\leq\delta$ (18)
Hence when the field is defined by $\Psi_{\text{att}}$ we have $W_{2}(P_{eq},Z_{subset})\leq\delta$.
Case 2: Let the field be defined by $\Psi_{\text{att}}^{*}$. The proxies $\textbf{p}_{k},k\in{1\ldots m}$ migrate to the nearest minimum in the potential field denoted by $z_{min,f(k)}$. Using Proposition 1 proved before, we know that all minima satisfy $\min_{j}\|z_{min,f(k)}-z_{j}\|>\delta$ for all $j$ except at most one $j=j^{\prime}$ for which let $k=k^{\prime}$.
First, we prove the corollary for the case if there exists such a $j=j^{\prime}$. Let $\textbf{z}_{g(k)},k=\{1\ldots m\}$ represent the ordered subset of data embeddings that minimize the $W_{2}$ distance metric with the proxies $\textbf{p}_{k}\in P_{eq}$. From proposition 1, we know that:
$\|z_{min,f(k)}-z_{j^{{}^{\prime}}}\|_{2}\leq\delta$
we also have:
$\|z_{g(k)}-z_{j^{{}^{\prime}}}\|_{2}\geq\min_{i,j}\|z_{i}-z_{j}\|_{2}\geq 2\delta\left(1+\dfrac{1}{m}\right)$
Using the triangle inequality on the above 2 inequalities and substituting, we get:
$\|z_{min,f(k)}-z_{g(k)}\|_{2}$ $\geq\|z_{g(k)}-z_{j^{{}^{\prime}}}\|_{2}-\|z_{min,f(k)}-z_{j^{{}^{\prime}}}\|_{2}$
$\|z_{min,f(k)}-z_{g(k)}\|_{2}$ $\geq 2\delta\left(1+\dfrac{1}{m}\right)-\delta$
$\|z_{min,f(k)}-z_{g(k)}\|_{2}$ $\geq\delta+\dfrac{2\delta}{m}$ (21)
The $W_{2}^{*}$ distance is given by:
$W_{2}^{*}(P_{eq},Z_{subset})=\inf_{\pi}\left(\dfrac{1}{m}\sum_{k=1}^{m}\|\textbf{p}_{k}-\textbf{z}_{\pi(k)}\|_{2}\right)$
here the infimum is over all permutations $\pi$ of k elements
$W_{2}^{*}(P_{eq},Z_{subset})=\left(\dfrac{1}{m}\sum_{k=1}^{m}\|z_{min,f(k)}-z_{g(k)}\|_{2}\right)$
$W_{2}^{*}(P_{eq},Z_{subset})=\left(\dfrac{1}{m}\sum_{k=1,k\neq k^{{}^{\prime}}}^{m}\|z_{min,f(k)}-z_{g(k)}\|_{2}\right)+\dfrac{1}{m}\|z_{min,k^{{}^{\prime}}}-z_{j^{\prime}}\|_{2}$
Using equation 21 and propositon 1
$W_{2}^{*}(P_{eq},Z_{subset})\geq\dfrac{1}{m}\times\left((m-1)\delta+\dfrac{2(m-1)\delta}{m}\right)$
$W_{2}^{*}(P_{eq},Z_{subset})>\delta$ (22)
Hence the $W_{2}$ distance between $P_{eq}$ and $Z_{subset}$ when the field is given by $\Psi_{att}$ is $W_{2}(P_{eq},Z_{subset})\leq\delta$ while their distance $W_{2}^{*}(P_{eq},Z_{subset})$ when the field is $>\delta$ as proved above. So $W_{2}^{*}(P_{eq},Z_{subset})>W_{2}(P_{eq},Z_{subset})$. hence, proved.
Appendix C Additional Experimental Results
In this section of the supplement, we provide additional experiments that we could not fit in the space available in the main paper. These empirical studies further validate the effectiveness of our method.
C.1 Performance using Small Embedding size
Context: In Section 4.2, we presented results on image retrieval using an embedding space of dimension 512. However in certain settings, learning embeddings in a lower dimension space might be more useful, such as in settings where limited storage is available for storing image embeddings. While this lowers the image retrieval performance, as is to be expected, it allows for a trade-off between available memory/compute resources and the accuracy of retrieval. Hence, we compare the performance of our method in learning a lower dimensional embedding space with recent state-of-the-art baselines.
Experiment: We train a ResNet-50 network with its embedding size set to 64 (a commonly used setting) on the Cars-196 [24], CUB-200-2011[47] and SOP [41] datasets.
Results: As seen in Tables 4 and 5, we observe that our method is able to outperform all other methods at this task; specifically, it outperforms strong Proxy-based baselines like ProxyAnchor[21] and ProxyGML[61] by more than 4.5 %, 2.2 % and 2.6% in the Recall@1 (R@1) metric on the Cars-196, CUB-200 and SOP datasets, respectively. It also outperforms the current state-of-the-art, the graph-based HIST[28] by 3% and 1.4% in terms of R@1 on the Cars-196 and CUB-200 datasets. This shows the strength of our method in learning a low-dimensional semantic representation space.
Benchmarks $\rightarrow$ CUB-200-2011 Cars-196 Methods $\downarrow$ R@1 R@2 R@4 R@8 R@1 R@2 R@4 R@8 MultiSimilarity [50] 57.4 69.8 80.0 87.8 77.3 85.3 90.5 94.2 SemiHard [38] 42.6 55.0 66.4 - 51.5 63.8 73.5 - LiftedStruct [41] 43.6 56.6 68.6 79.6 53.0 65.7 76.0 84.3 N-Pair [40] 51.0 63.3 74.3 83.2 71.1 79.7 86.5 91.6 ProxyNCA [33] 49.2 61.9 67.9 72.4 73.2 82.4 86.4 88.7 SoftTriple [35] 60.1 71.9 81.2 88.5 78.6 86.6 91.8 95.4 Clustering [42] 48.2 61.4 71.8 81.9 58.1 70.6 80.3 87.8 ProxyAnchor [21] 61.7 73.0 81.8 88.8 78.8 87.0 92.2 95.5 ProxyGML [61] 59.4 70.1 80.4 - 78.9 87.5 91.9 - HIST [28] 62.5 73.6 83.0 89.6 80.4 87.6 92.4 95.4 Ours 63.9 $\pm$ 0.3 74.7 $\pm$ 0.2 83.5 $\pm$ 0.1 90.1 $\pm$ 0.1 83.4 $\pm$ 0.2 89.9 $\pm$ 0.2 94.2 $\pm$ 0.1 97.1 $\pm$ 0.1
[…]
Benchmarks $\rightarrow$ SOP
Methods $\downarrow$ R@1 R@10 R@100 R@1000
MultiSimilarity [50] 74.1 87.8 94.7 98.2
LiftedStruct [41] 62.5 80.8 91.9 -
N-Pair [40] 67.7 83.8 93.0 97.8
ProxyNCA [33] 73.7 - - -
SoftTriple [35] 76.3 89.1 95.3 -
Clustering [42] 67.0 83.7 93.2 -
ProxyAnchor [21] 76.5 89.0 95.1 98.2
ProxyGML [61] 76.2 89.4 95.4 -
HIST [28] 78.9 90.5 95.8 98.5
Ours 79.5 $\pm$ 0.2 90.8 $\pm$ 0.1 96.3 $\pm$ 0.1 98.6 $\pm$ 0.1
[…]
C.2 Evaluation under MLRC Protocol
[…]
Table 6: Comparison of the Precision@1, R-Precision (RP) and the Mean Average Precision @ R (MAP@R) as defined in [34] achieved by our method on the Cars-196 dataset with state-of- the-art baselines under MLRC[34] settings.
Embedding type $\rightarrow$ Concatenated (512-dim) Separated (128-dim) Methods $\downarrow$ P@ 1 RP MAP@R P@ 1 RP MAP@R Contrastive [8] $68.1\pm 0.3$ $37.2\pm 0.3$ $26.5\pm 0.3$ $59.7\pm 0.4$ $32.0\pm 0.3$ $21.2\pm 0.3$ Triplet [51] $64.2\pm 0.3$ $34.6\pm 0.2$ $23.7\pm 0.2$ $55.8\pm 0.3$ $29.6\pm 0.2$ $18.8\pm 0.2$ N-Pair [40] $66.6\pm 0.3$ $36.0\pm 0.2$ $25.1\pm 0.2$ $58.1\pm 0.2$ $30.8\pm 0.2$ $19.9\pm 0.2$ ProxyNCA [33] $65.7\pm 0.4$ $35.1\pm 0.3$ $24.2\pm 0.3$ $57.9\pm 0.3$ $30.2\pm 0.2$ $19.3\pm 0.2$ Margin [52] $63.6\pm 0.5$ $33.9\pm 0.3$ $23.1\pm 0.3$ $54.8\pm 0.3$ $28.9\pm 0.2$ $18.1\pm 0.2$ Margin/class [52] $64.4\pm 0.2$ $34.6\pm 0.2$ $23.7\pm 0.2$ $55.6\pm 0.2$ $29.3\pm 0.2$ $18.5\pm 0.1$ N. Softmax [58] $65.6\pm 0.3$ $36.0\pm 0.2$ $25.3\pm 0.1$ $58.8\pm 0.2$ $31.8\pm 0.1$ $21.0\pm 0.1$ CosFace [49] $67.3\pm 0.3$ $37.5\pm 0.2$ $26.7\pm 0.2$ $59.6\pm 0.4$ $32.0\pm 0.2$ $21.2\pm 0.2$ ArcFace [10] $67.5\pm 0.3$ $37.3\pm 0.2$ $26.5\pm 0.2$ $60.2\pm 0.3$ $32.4\pm 0.2$ $21.5\pm 0.2$ FastAP [4] $63.2\pm 0.3$ $34.2\pm 0.2$ $23.5\pm 0.2$ $55.6\pm 0.3$ $29.7\pm 0.2$ $19.1\pm 0.2$ SNR [56] $66.4\pm 0.6$ $36.6\pm 0.3$ $25.8\pm 0.4$ $58.1\pm 0.4$ $31.2\pm 0.3$ $20.4\pm 0.3$ MultiSimilarity [50] $65.0\pm 0.3$ $35.4\pm 0.1$ $24.7\pm 0.1$ $57.6\pm 0.2$ $30.8\pm 0.1$ $20.2\pm 0.1$ MS+Miner [50] $67.7\pm 0.2$ $37.3\pm 0.2$ $26.5\pm 0.2$ $59.4\pm 0.3$ $31.9\pm 0.1$ $21.0\pm 0.1$ SoftTriple [35] $67.3\pm 0.4$ $37.3\pm 0.2$ $26.5\pm 0.2$ $59.9\pm 0.3$ $32.1\pm 0.1$ $21.3\pm 0.1$ ProxyAnchor $65.2\pm 0.2$ $36.0\pm 0.2$ $25.3\pm 0.1$ $56.6\pm 0.1$ $30.5\pm 0.1$ $19.8\pm 0.2$ HIST [28] $69.6\pm 0.3$ $38.8\pm 0.1$ $28.2\pm 0.1$ $61.3\pm 0.2$ $33.1\pm 0.2$ $22.3\pm 0.1$ Ours $\textbf{70.1}\pm\textbf{0.2}$ $\textbf{39.9}\pm\textbf{0.1}$ $\textbf{29.4}\pm\textbf{0.3}$ $\textbf{62.4}\pm\textbf{0.1}$ $\textbf{33.8}\pm\textbf{0.2}$ $\textbf{23.1}\pm\textbf{0.3}$
Table 7: Comparison of the Precision@1, R-Precision (RP) and the Mean Average Precision @ R (MAP@R) as defined in [34] achieved by our method on the CUB-200-2011 dataset with state-of-the-art baselines under MLRC[34] settings.
Context: In Section4.2, we compared the image retrieval performance of our method with other recent techniques using standard evaluation settings used in [21, 28, 60, 29]. Recently, some studies [34, 36] have pointed to flaws in these settings, including the lack of a standardized backbone architecture, weakness of the metrics used, and the lack of a standardized validation subset. Though we address some of these flaws by comparing against methods using the same experimental settings (backbone, embedding dimensions, image sizes) as described in Section 4.1, in this section we additionally evaluate our method under the constrained protocol proposed in [34]. The constrained protocol proposes using fixed optimization settings with no learning rate scheduling to train an Inception with BatchNorm architecture. It also introduces new, more informative metrics (the R-Precision and Mean Average Precision@R). Further details of the constrained protocol can be found in  [34].
Results: We evaluate the performance of our method using models trained under the constrained protocol on the Cars-196 [24] and CUB-200-2011 [47] datasets. As seen in Table 6, our method significantly outperforms all previous methods on all metrics. We significantly outperform the previous best pair-based method, the MultiSimilarity loss [50] by 7.4% and 4.8 % in terms of P@1 (128-dim embeddings) on the Cars-196 and CUB-200 datasets respectively. We also outperform the current state-of-the-art method [28], HIST by 1.9 % and 1.1 % in terms of P@1 (128-dim) on the Cars-196 and CUB-200 datasets respectively. We note that these gains are higher than the improvements made by the current state-of-the-art HIST (1.5% and 1.1 % ) over previous methods on these benchmarks. Our method also outperforms all previous methods in terms of the R-Precision and Mean Average Precision @ R metric, demonstrating the quality of the semantic metric learned by it.
Appendix D Visual Results
We present qualitative results for image retrieval by our method to evaluate the semantic similarity metric learned by it. Figure 4 displays 2 examples of query images from each of the 3 datasets, followed by 4 nearest images retrieved by our method, arranged in increasing order of distance. It can be seen that despite the large intra-class variation (pose, color) in the datasets, our method is able to effectively retrieve similar images.
Figure 4: Example image retrieved by our method for query images from (a) Cars-196 (b) CUB-200-2011 and (c) SOP test datasets, in increasing order of distance from the query. Correct retrievals have a green border, while incorrect ones have a red one.
Figure 5 displays a t-sne visualization of the embedding space learnt by our method on the CUB-200-2011 dataset. it can be seen that images closer together share more semantic characteristics than those that are far apart.
Figure 5: A t-sne visualization of a semantic representation space learnt by our method on the CUB-200 dataset